# Extremal metrics involving scalar curvature

### v2

@@ 4 Proofs of Main Theorems
and $R_{g}\equiv\mathrm{const}$, $\overset{\circ}{\mathrm{Ric}_{g}}$ is divergence-free. Since $\overset{\circ}{\mathrm{Ric}_{g}}$ is also trace-free, hence it is a tt-tensor. We are going to take this tt-tensor $\overset{\circ}{\mathrm{Ric}_{g}}$ as the variation $h$ in the following proof of our main theorems.
@@
From the formula (15), we have for $10<s<<$ 1,
@@
The first one is known as the first variation of the scalar curvature functional along the Ricci flow (see [1, Corollary 4.20] for example). Here, $\overline{\lambda}_{Ric}(x)~{}(x\in M)$ is the largest eigenvalue of $\mathrm{Ric}_{g}(x)$ on $T_{x}M$, and $Q_{s}(\nabla^{g}\mathrm{Ric}_{g})=O(s^{2})\cdot Q(\nabla^{g}\mathrm{Ric}_{g})$, where $Q(\nabla^{g}\mathrm{Ric}_{g})$ is a contraction of three copies of $g_{s}^{-1}$ and two copies of $\nabla^{g}\mathrm{Ric}_{g}$.
@@
Let $\lambda_{Ric,1}(x)\leq\cdots\leq\lambda_{Ric,n-1}(x)\leq\lambda_{Ric,n}(x)$ be the eigenvalues of $\mathrm{Ric}_{g}(x)$ on $T_{x}M~{}(x\in M)$. As in the proof of Theorem 1.3, we have that for all sufficiently small $10<|s|<<$ 1 (so that $(1+(R_{g}/n)s)g-s\mathrm{Ric}_{g}$ is positive definite on $M$),
@@
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}-\frac{R_{g}}{\sqrt{\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n}\right)\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n-1}\right)}}+O(s^{2})\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,n-1}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\left(\frac{s\lambda_{Ric,n-1}}{\left(1-s\lambda_{Ric,n-1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}~{}-\frac{sR_{g}}{\left(1-s\lambda_{Ric,n-1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,n-1}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
@@
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}-\frac{R_{g}}{\sqrt{\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n}\right)\left(1+\frac{R_{g}}{n}s-s\lambda_{Ric,n-1}\right)}}+O(s^{2})\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,n}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\left(\frac{s\lambda_{Ric,n}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~{}~{}~{}~{}-\frac{sR_{g}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,n}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
@@
For the case of $s<0$, one can prove the desired assertion by replacing $\lambda_{Ric,n},\lambda_{Ric,n-1}$ with $\lambda_{Ric,2},\lambda_{Ric,1}$ respectively for each case: $R_{g}\leq$ and $\geq 0$ respectively. (Of course, the second inequality sign should be the opposite for each estimate in this case.) ∎
@@
Proof of Theorem 1.7.
Let $h$ be a tt-tensor on $(M,g)$. Then, as discussed in [16, Remark 1.6], one can find a curve $g_{s}~{}(|s|<<1)$ lying in the space of all Riemannian metrics of constant scalar curvature with unit volume such that $\frac{d}{ds}g_{s}|_{s=0}=h$. By the Taylor expansion, we have that
$R_{g_{s}}=R_{g}+DR|_{g}(h)\cdot s+\frac{1}{2}\left(D^{2}R|_{g}(h,h)+DR|_{g}\left(h_{2}\right)\right)\cdot s^{2}+O(s^{3})$
on $M$. Here, $h_{2}:=\left.\frac{d^{2}}{d^{2}s}g_{s}\right|_{s=0}$. Since $g$ is an Einstein metric and $h$ is a tt-tensor, $DR|_{g}(h)=0$. Hence, for any $\varepsilon>0$ there is a small $s>0$ such that
$\begin{split}R_{g_{s}}-R_{g}\cdot&\|g\|^{2}_{1,g_{s}}\geq R_{g}+\frac{1}{2}\left(D^{2}R|_{g}(h,h)+DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})-\frac{R_{g}}{1+s(h_{min}\pm\varepsilon)}\\ &=R_{g}+\frac{1}{2}\left(D^{2}R|_{g}(h,h)+DR|_{g}(h_{2})\right)\cdot s^{2}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\left(1-s(h_{min}\pm\varepsilon)+s^{2}(h_{min}\pm\varepsilon)^{2}\right)+O(s^{3})\\ &=sR_{g}(h_{min}\pm\varepsilon)-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})+R_{g}(h_{min}\pm\varepsilon)^{2}\right)s^{2}+O(s^{3}),\end{split}$
Here,
±$\pm\varepsilon:=\begin{cases}+\varepsilon&\mathrm{if}~{}R_{g}\leq 0~{}\mathrm{and}\\ -\varepsilon&\mathrm{if}~{}R_{g}\geq 0,\end{cases}$
$h_{min}(x):=\sup\{C\in\mathbb{R}\mid h_{x}(v,v)\geq C\cdot g_{x}(v,v)~{}\mathrm{for~{}all}~{}v\in T_{x}M\}\leq 0$ since $\mathrm{tr}_{g}h=0$, and we have used
$D^{2}R|_{g}(h,h)=-\frac{1}{2}\langle\Delta_{E}h,h\rangle_{L^{2}}$
since $g_{s}$ lies in the space of all Riemannian metrics of constant scalar curvature (see [16, Remark 1.6]). Therefore, if $R_{g}=0$ and $DR_{g}(h_{2})=0$, then $R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}$ on $M$ for all sufficiently small $s>0$. If $R_{g}<0$ and $\max_{M}h_{min}<0$, then one can take $s>0$ small enough so that $h_{min}+\varepsilon<0$ on $M$. Therefore, $R_{g_{s}}>R_{g}\cdot\|g\|^{2}_{1,g_{s}}$ on $M$ for all sufficiently small $s>0$. Note that since $\mathrm{tr}_{g}h=0$, $\max_{M}h_{min}<0$ is equivalent to that $h$ is non-vanishing on $M$. ∎
Proof of Theorem 1.8.
As in the proof of Theorem 1.6, for any $\varepsilon>0$ there is a small $s>0$ such that
$\begin{split}&R_{g_{s}}-R_{g}\cdot\|g\|_{2,g_{s}}\geq R_{g}-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\cdot\begin{cases}\frac{1}{\sqrt{(1+s(\lambda^{h}_{1}+\varepsilon))(1+s(\lambda^{h}_{2}+\varepsilon))}}&\mathrm{if}~{}R_{g}\leq 0\\ \frac{1}{\sqrt{(1+s(\lambda^{h}_{1}-\varepsilon))(1+s(\lambda^{h}_{2}-\varepsilon))}}&\mathrm{if}~{}R_{g}\geq 0\end{cases}\\ &\geq R_{g}-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})+\begin{cases}\frac{-R_{g}}{1+s(\lambda^{h}_{2}+\varepsilon)}&\mathrm{if}~{}R_{g}\leq 0\\ \frac{-R_{g}}{1+s(\lambda^{h}_{1}-\varepsilon)}&\mathrm{if}~{}R_{g}\geq 0\end{cases}\\ &=R_{g}-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})\right)\cdot s^{2}+O(s^{3})\\ &~{}~{}~{}~{}~{}~{}~{}~{}-R_{g}\cdot\begin{cases}(1-s(\lambda^{h}_{2}+\varepsilon)+s^{2}(\lambda^{h}_{2}+\varepsilon)^{2})&\mathrm{if}~{}R_{g}\leq 0\\ (1-s(\lambda^{h}_{1}-\varepsilon)+s^{2}(\lambda^{h}_{1}-\varepsilon)^{2})&\mathrm{if}~{}R_{g}\geq 0\end{cases}\\ &=\begin{cases}sR_{g}(\lambda^{h}_{2}+\varepsilon)-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})+R_{g}(\lambda^{h}_{2}+\varepsilon)^{2}\right)s^{2}+O(s^{3})&\mathrm{if}~{}R_{g}\leq 0\\ sR_{g}(\lambda^{h}_{1}-\varepsilon)-\left(\frac{1}{4}\langle\Delta_{E}h,h\rangle_{L^{2}}-\frac{1}{2}DR|_{g}(h_{2})+R_{g}(\lambda^{h}_{1}-\varepsilon)^{2}\right)s^{2}+O(s^{3})&\mathrm{if}~{}R_{g}\geq 0.\end{cases}\end{split}$
Similarly, if $s<0$ with sufficiently small $|s|$, we have the same estimate by replacing $\lambda^{h}_{1},\varepsilon$ with $\lambda^{h}_{2},-\varepsilon$ respectively. Note that $\lambda^{h}_{1}\leq 0$ since $\mathrm{tr}_{g}h=0$, hence the desired assertions follow from these estimates. ∎
@@ 5.1 Examples on Lie groups
  • A left invariant metric on $SU(2)$ (which is homeomorphic to the unit 3-sphere $\mathbb{S}^{3}$), whose signature of Ricci quadratic form is (+,,)(+,-,-)( + , - , - ) (see also [12, Example 2]). More specifically, we consider the Berger sphere:
@@
    Hence, if $40<\varepsilon<$ 4, then $g_{s,t}$ has a positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6. Moreover, from [12, Example 2], if $\varepsilon\leq 2$ (resp. $\varepsilon<2$), then $g_{s,t}$ ia a (resp. unique) Yamabe metric with positive scalar curvature. On the other hand, if $\varepsilon\leq 0$, then $g_{s,t}$ has a non-positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4 and 1.5 but does not one of Theorem 1.6.
@@
    whose signature of Ricci quadratic form is (+,,)(+,-,-)( + , - , - ).
  • A left invariant metric on $SL(2,\mathbb{R})$ or $E(1,1)$, whose signature of Ricci quadratic form can be either (+,,)(+,-,-)( + , - , - ) or $(0,0,-)$ depending on the choice of the left invariant metric.
@@
5.2 Examples on the total spaces of Riemannian submersions with totally geodesic fibres
  • ([25, Section 5]) Let $\pi:S^{4n+3}\rightarrow\mathbb{H}P^{n}$ be the Hopf fibration whose fibres are the standard unit $3$-sphere $\mathbb{S}^{3}$. We denote the scalar curvatures of the fibres, the base space and the total space as $R^{F},R^{B}$ and $R^{M}$, respectively. Then,
@@ 5.2 Examples on the total spaces of Riemannian submersions with totally geodesic fibres
    Therefore, $(\mathbb{S}^{4n+3},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $10<t<<$ 1 and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $10<t<<$ 1 and $t>>1$.
  • ([25, Section 5]) Consider the fibration $\mathbb{C}P^{2n+1}\rightarrow\mathbb{H}P^{n}$ whose fibres are $S^{2}(4)$ with constant sectional curvature $4$. Then,
@@
Therefore, $(\mathbb{C}P^{2n+1},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $10<t<<$ 1 and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $10<t<<$ 1 and $t>>1$.

### v3

@@ 4 Proofs of Main Theorems
and $R_{g}\equiv\mathrm{const}$, $\overset{\circ}{\mathrm{Ric}_{g}}$ is divergence-free. Since $\overset{\circ}{\mathrm{Ric}_{g}}$ is also trace-free, hence it is a tt-tensor. We are going to take this tensor $\overset{\circ}{\mathrm{Ric}_{g}}$ as the variation $h$ in the following proof of our main theorems.
@@
From the formula (13), we have for $0<s<<1$,
@@
The first one is known as the first variation of the scalar curvature functional along the Ricci flow (see [2, Corollary 4.20] for example). Here, $\overline{\lambda}_{Ric}(x)~(x\in M)$ is the largest eigenvalue of $\mathrm{Ric}_{g}(x)$ on $T_{x}M$, and $Q_{s}(\nabla^{g}\mathrm{Ric}_{g})=O(s^{2})\cdot Q(\nabla^{g}\mathrm{Ric}_{g})$, where $Q(\nabla^{g}\mathrm{Ric}_{g})$ is a contraction of three copies of $g_{s}^{-1}$ and two copies of $\nabla^{g}\mathrm{Ric}_{g}$.
@@
Let $\lambda_{Ric,1}(x)\leq\cdots\leq\lambda_{Ric,n-1}(x)\leq\lambda_{Ric,n}(x)$ be the eigenvalues of $\mathrm{Ric}_{g}(x)$ on $T_{x}M~(x\in M)$. As in the proof of Theorem 1.3, we have that for all sufficiently small $0<|s|<<1$ (so that $(1+(R_{g}/n)s)g-s\mathrm{Ric}_{g}$ is positive definite on $M$),
@@
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,1}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~~~~~~~~~~-R_{g}\left(\frac{s\lambda_{Ric,1}}{\left(1-s\lambda_{Ric,1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~~~~-\frac{sR_{g}}{\left(1-s\lambda_{Ric,1}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,1}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
@@
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &\geq s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}+\frac{R_{g}}{1+s\frac{R_{g}}{n}}-\frac{R_{g}}{1-s\lambda_{Ric,n}+\frac{R_{g}}{n}s}+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}+s\frac{R^{2}_{g}}{n}\\ &~~~~~~~~~~-R_{g}\left(\frac{s\lambda_{Ric,n}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\right)+O(s^{2})\\ &=s\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{g}+\frac{3n-2}{2n}\Delta_{g}R_{g}\\ &~~~~-\frac{sR_{g}}{\left(1-s\lambda_{Ric,n}+s\frac{R_{g}}{n}\right)\left(1+s\frac{R_{g}}{n}\right)}\left(\lambda_{Ric,n}-\frac{R_{g}}{n}+O(s)\right)+O(s^{2}).\end{split}$
@@
For the case of $s<0$, one can prove the desired assertion by replacing $\lambda_{Ric,1},\lambda_{Ric,n}$ with $\lambda_{Ric,n},\lambda_{Ric,1}$ respectively for each case: $R_{g}\leq$ and $\geq 0$ respectively. (Of course, the second inequality sign should be the opposite for each estimate in this case.) ∎
@@ 5 Examples
5.1 On the product of two Einstein spaces
Let $n,m\geq 2$ and $(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$, $(\mathbb{H}^{m},g_{\mathbb{H}^{m}})$ be space forms with sectional curvature $1$ and 1-1 respectively. Consider the metric of the form
$g_{\lambda}:=g_{\mathbb{S}^{n}}+\lambda g_{\mathbb{H}^{m}},$
where $\lambda\geq 1$ is a scaling constant of the hyperbolic space. Then, for each point $x\in\mathbb{S}^{n}\times\mathbb{H}^{m}$,
$\mathrm{Ric}(g_{\lambda})=\begin{cases}(n-1)g_{\mathbb{S}^{n}}&\mathrm{on}~T_{\mathbf{p}_{1}x}\mathbb{S}^{n}\subset T_{x}(\mathbb{S}^{n}\times\mathbb{H}^{m}),\\ -(m-1)g_{\mathbb{H}^{m}}&\mathrm{on}~T_{\mathbf{p}_{2}x}\mathbb{H}^{m}\subset T_{x}(\mathbb{S}^{n}\times\mathbb{H}^{m}),\end{cases}$
and hence
$R(g_{\lambda})=n(n-1)-\frac{m(m-1)}{\lambda}.$
Here, $\mathbf{p}_{i}~(i=1,2)$ denotes the natural projection from respectively $\mathbb{S}^{n}$ and $\mathbb{H}^{m}$ to $\mathbb{S}^{n}\times\mathbb{H}^{m}$.
So, if we put $n=m$, then one can check that $R(g_{\lambda})\geq 0$ and $g_{\lambda}$ satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6 for all $\lambda\geq 1$. Moreover, if we put $n=m+1$, one can also check that $R(g_{1})\geq 0$ and $g_{1}$ satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6. On the other hand, we can directly check that $g_{1}$ is none of type I, II, III or IV scalar curvature rigid in the sense of Listing. In fact, the family $\{g_{\lambda}\}_{\lambda\geq 1}$ gives a deformation which suggests that $g_{1}$ is not rigid in each sense. In particular, for every dimension $n\geq 4$, there is an $n$-dimensional manifold on which there is a non-rigid metric $g$ in each sense.
@@ 5.2 Examples on Lie groups
  • A left invariant metric on $SU(2)$ (which is homeomorphic to the unit 3-sphere $\mathbb{S}^{3}$), whose signature of Ricci quadratic form is (+,,)(+,-,-) (see also [14, Example 2]). More specifically, we consider the Berger sphere:
@@
    Hence, if $0<\varepsilon<4$, then $g_{s,t}$ has a positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6. Moreover, from [14, Example 2], if $\varepsilon\leq 2$ (resp. $\varepsilon<2$), then $g_{s,t}$ is a (resp. unique) Yamabe metric with positive scalar curvature. On the other hand, if $\varepsilon\leq 0$, then $g_{s,t}$ has a non-positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4 and 1.5 but does not one of Theorem 1.6.
@@
    whose signature of Ricci quadratic form is (+,,)(+,-,-).
  • A left invariant metric on $SL(2,\mathbb{R})$ or $E(1,1)$, whose signature of Ricci quadratic form can be either (+,,)(+,-,-) or (0,0,)($0,$-) depending on the choice of the left invariant metric.
@@
5.3 Examples on the total spaces of Riemannian submersions with totally geodesic fibers
  • ([26, Section 5]) Let $\pi:S^{4n+3}\rightarrow\mathbb{H}P^{n}$ be the Hopf fibration whose fibers are the standard unit $3$-sphere $\mathbb{S}^{3}$. We denote the scalar curvatures of the fibers, the base space and the total space by $R^{F},R^{B}$ and $R^{M}$, respectively. Then,
@@ 5.3 Examples on the total spaces of Riemannian submersions with totally geodesic fibers
    Therefore, $(\mathbb{S}^{4n+3},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $0<t<<1$ and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $0<t<<1$ and $t>>1$.
  • ([26, Section 5]) Consider the fibration $\mathbb{C}P^{2n+1}\rightarrow\mathbb{H}P^{n}$ whose fibers are $S^{2}(4)$ with constant sectional curvature $4$. Then
@@
Therefore, $(\mathbb{C}P^{2n+1},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $0<t<<1$ and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $0<t<<1$ and $t>>1$.
Question 5.1.
Are there any manifolds of non-positive Yamabe invariants that satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6?